OpenStudy (anonymous):

2sinXcosX = cosX solve for x, how can i do it

OpenStudy (anonymous):

i can only find 2 solutions: pi/6 and 5pi/6, but the solution says there are 4 solutions, how can i find the other 2?

OpenStudy (anonymous):

2sinxcosx=cosx 2sinxcosx - cosx =0 cosx (2sinx - 1) = 0 take cos x=0 solve for x ... or 2sinx - 1 = 0 --> sinx = 1/2 solve x

OpenStudy (anonymous):

Do u know cosx=0, for x=....? and sinx=1/2, for x = ...?

OpenStudy (anonymous):

@tanjung oh i see.. thanks u are a beast!

OpenStudy (anonymous):

cosx=0, for x=pi/2 and 3pi/2 cosx=1/2, for x=pi/3 and 5pi/3 there are 4 solution